OpenStudy (anonymous):

Y=4x and y+x=5 A. (-4,1) B. (1,4) C. (-3,2) D. (2,3)

OpenStudy (er.mohd.amir):

put y from first to second that is 4x+x=5

OpenStudy (er.mohd.amir):

5x=5 x=5/5 x=1 y=4x=4*1=4